OpenStudy (anonymous):

what is the formula for the surface area of a cylinder in terms of the radius? (i.e. solve for 'r')

OpenStudy (dangerousjesse):

It's \(A=2πrh+2πr2\)

OpenStudy (anonymous):

thank you but like i need the equation to equal r

OpenStudy (nikato):

I don't think that's what the question is asking. When it says in terms of r, it just means to let r represent the radius It doesn't mean to make it r=...

OpenStudy (nikato):

So that formula^ from Jesse is the correct answer

OpenStudy (anonymous):

the actual problem says surface area of cylinder 54(pi) and the height is 6cm, solve in terms of radius, @nikato

OpenStudy (wolf1728):

What is the formula from Jesse? So far I have Area / (2 PI h + 2 PI r) = r which still leaves an r on both sides

OpenStudy (anonymous):

jesse just put the formula for the surface area of a cylinder.. @wolf1728

OpenStudy (nikato):

@kshirley oh, you should have said that earlier so you have to plug in 54pi for A and 6 for h in the formula

OpenStudy (nikato):

so \[54\pi= 2\pi r(6) + 2\pi r ^{2}\]

OpenStudy (wolf1728):

I thought you need it solved for 'r' The best I could do is: h*r + r² = Area / (2 PI)

OpenStudy (wolf1728):

Here's how to solve it: http://mathcentral.uregina.ca/QQ/database/QQ.09.10/h/jazmin1.html Basically, it requires the quadratic formula where a=1 b=height and c= -1*(Area/(2*PI)) Let's suppose Area = 150.7964473723 sq inches and height = 5 inches What is the radius? a=1 b=5 c=-24 x = (-5 +- sqrt(25 - 4*-24)) / 2*1 x = -5 +- sqrt (25 +96) / 2 x = -5 +(11)/ 2 x = 6/2 radius = 3 which it does !!!

OpenStudy (wolf1728):

Thanks for the medal!!!

OpenStudy (anonymous):

thank you SOOO much!!!

OpenStudy (wolf1728):

u r welcome :-)